You want to start an organic garlic farm. The farm costs $190,000, to be paid in full immediately. Year 1 cash flows will be $25,000, and grow at 5% a year into year 5 when you decide to sell the farm at the end of the year for $300,000 (in year 5 you get cash flows from selling both the garlic and the farm). Assuming these estimates are all correct, what is the IRR of the garlic farm investment? Round to the tenth of a percent (e.g. 5.6%=5.6). ​[Hint: You'll want to solve this in Excel using the IRR function or Goal Seek. Make sure the IRR cell is set to show decimals.]
Answer #1

Let the cash flows for Year n be denoted by CFn

Given,

CF0 = -190000

CF1 = 25000

CF2 = 25000(1+0.05) = 26250

CF3 = 26250(1+0.05) = 27562.5

CF4 = 27562.5(1+0.05) = 28940.63

CF5 = 28940.63(1+0.05) + 300000 = 330387.66

Using excel IRR function to compute IRR

B8 V x B fx =IRR(B2:17) C D 1 Year Ham to do بر با بيه 1 A Cash Flow 0 -190000.00 25000.00 2 26250.00 3 27562.50 4 28940.63 3

Hence, IRR = 21.78%